# Do More Suspicious Transaction Reports Lead to More Convictions for Money Laundering?

### v1

@@ Abstract
Our results have important implications for international anti-money laundering efforts and policies. At best, the number of suspicious transaction reports and convictions for money laundering in a country follow a sub-linear power law. At worst, there is no or only a very weak (and insignificant) relationship between the number of suspicious transaction reports and convictions for money laundering in a country. Either way, our results suggest that financial regulators and supervisors should focus less on the number of suspicious transaction reports filed in a given country. Instead, they should focus more on how reports are utilized differently across countries and over time.
@@ 1 Introduction
A meta-analysis published by the United Nations Office on Drugs and Crime [61] suggests that money laundering amounts to 2.7% of the world’s combined gross domestic product (GDP). At the same time, scholars have long questioned the effectiveness of global anti-money laundering (AML) efforts [63, 58, 56, 62]. Such efforts are guided by the Financial Action Task Force (FATF), an intergovernmental organization founded in 1989. Among other things, FATF recommends that entities such as banks should be required to report suspicious transactions to national financial intelligence units [50]. The reports are known as suspicious transaction or activity reports (we use the former term) and are intended to help authorities detect and prosecute money laundering. In this paper, we investigate the relationship between suspicious transaction reports (STRs) and convictions for money laundering in the European Union (EU). Following the EU’s 4th AML directive, EU countries must have “effective” AML systems and, among other things, require obligated entities to file STRs [45]. More or less implicitly, the EU directive and FATF’s AML recommendations thus assume that STRs are an effective tool to combat money laundering. This is also evident from the fact that STR counts are used as a measure of success in international AML evaluations [49, 48]. Our article explores this idea, asking if “more STRs lead to more convictions for money laundering?”
@@
Our results have important implications for international AML efforts and policies. At best, the number of STRs and convictions for money laundering in a country follow a sub-linear power law. At worst, there is no or only a very weak (and insignificant) relationship between the number of STRs and convictions for money laundering in a country. This would be a cause for serious concern about the effectiveness of current AML efforts and policies. Either way, our results suggest that financial regulators and supervisors should focus less on increasing the number of STRs filed in a given country. Instead, they should focus more on how STRs are utilized differently across countries and over time.
@@ 2 Related Literature
Scholars have long questioned the effectiveness of global AML efforts and policies [63, 58, 56, 62]. [62] calls AML “the world’s least effective policy experiment,” arguing that less than 0.1% of criminal proceeds are seized while AML costs vastly exceed assets recovered. The author partly ascribes the failure of the global AML system to an excessive focus on activity metrics (e.g., the number of STRs filed) over outcomes (e.g., the number of convictions achieved or reduced crime rates). Somewhat more diplomatic, [58] argue that evaluation of AML efforts is hindered by an absence of reliable data. Indeed, the authors note that AML evaluations and assessments (including those by FATF) mostly rely on rhetorical arguments and narrative judgments, providing little to no evidence on effectiveness. This fits with a broader pattern, noted by [57], whereby “governments and intergovernmental organizations have spent almost nothing on public research on these issues, while regular costly (…) international meetings service the global fight against this ill-understood phenomenon and to ‘evaluate’ these efforts.” In an extensive essay and review, [56] note that ”the universal message from the modest research that has been conducted is that the proportion of SARs [a version of STRs] in high-reporting jurisdictions that are actually seriously followed up is low, though the extent to which this is inherent or merely resource-constrained remains unclear.” In a broader perspective, [56] note that “how well the [AML] system works in suppressing crimes and preventing terrorist acts is entirely a matter of speculation.”
@@ 3.3 Estimates of the Shadow Economy from the International Monetary Fund
We stress that “the shadow economy” and “money laundering” are distinct concepts. In particular, the shadow economy is defined as economic activity hidden from authorities [55, 54, 53, 69]. This will largely include activities that, in and of themselves, are legal. Furthermore, such activities do not necessarily involve money laundering; just as all criminal activities do not require money laundering. [^4] Imagine, for example, a day laborer who is paid in cash. The laborer’s work may, in and of itself, be legal. If the laborer fails to declare their income, it constitutes tax evasion. However, if the laborer simply uses their cash salary (including what should have been paid in taxes) to pay for daily expenses such as groceries, they may not be involved in any kind of money laundering.
@@
We use the European Sourcebook of Crime and Criminal Justice Statistics [37, 36, 38] as our source of police and conviction statistics. [^5] We extract raw statistics from the Sourcebook’s databases, available on the Sourcebook’s website [37]. The Sourcebook collects crime and criminal justice statistics from national correspondents. The 6th edition of the Sourcebook, for example, contains statistics from the Danish Ministry of Justice, the Latvian Statistical Bureau, and numerous academics from across Europe. Notably, the Sourcebook aims to collect standardized statistics; a prerequisite for comparing statistics across countries. [^6] As an alternative data source, we considered using mutual evaluation reports from FATF. However, statistics in these are not standardized. For example, [47] combines statistics on money laundering and stolen goods. While the Sourcebook aims to collect standardized statistics, these may deviate. In this work, we simply use the statistics as given in the Sourcebook. The Sourcebook explicitly defines money laundering as “specific financial transactions to conceal the identity, source, and/or destination of money or non-monetary property deriving from criminal activities.” We collect the following variables from the Sourcebook:
@@
Figures 2 and 3 illustrate the number of observations of $CON_{(c,t)}$ per country and year in our dataset. As can be seen, there are some countries for which we have no observations. Still, our dataset covers North, South, Eastern, and Western Europe. As can also be seen, the number of observations increases towards the end of our data collection period. We believe this is due to improvements in the way data was collected in the Sourcebook; not interaction effects between STRs and convictions.
@@ 4 Methodology
denote the average number of convictions in country $c$ per year up to and including $H$ years after $t$. We measure $CON_{(c,t+j)}$ and $CON_{(c,t+1:H)}$ per 100,000 capita using $POP_{(c,t)}$, i.e., the population size of country $c$ in year tt. [^11] This is based on the idea that $CON_{(c,t+j)}$ and $CON_{(c,t+1:H)}$ are determined by $STR_{(c,t)}$, i.e., the number of STRs in year $t$, not years t+1,,t+Ht+1,...,t+H. For robustness, we also tried measuring $CON_{(c,t+j)}$ and $CON_{(c,t+1:H)}$ using population counts in year $t+j$ and the average population count over years t+1,,t+Ht+1,...,t+H. This does not materially change our conclusions. We consider and model:
@@ 4.1 Pooled Regression
where $A\in\mathbb{R}_{>0}$ measures a “scale of production” and $\beta\in\mathbb{R}$ measures how convictions respond to a change in STRs. To fit the model, we use a log-transformation, yielding an ordinary least squares regression
@@
Motivated by the observation that we have per capita ratios on both sides of equation (3) (note that both $CON_{(c,t+j)}$ and $STR_{(c,t)}$ are measured per 100,000 capita) and that a country’s population size might have a stand-alone effect on conviction rates, we include $log_{10}(POP_{(c,t)})$ as a control variable in the model in (3). [^14] Algebraically, $\log_{10}(y/n)=\alpha+\beta\,\log_{10}(x/n)+u\iff\log_{10}y=\alpha+\beta\,\log_{10}(x)+(1-\beta)\,\log_{10}(n)+u$. Thus, omitting $\log_{10}(n)$ (or, in our case, $\log_{10}(POP_{(c,t)})$) as a control variable imposes a (hidden) restriction that its coefficient must equal $1-\beta$. Note, also, that measuring variables per (100,000) capita shifts the intercept of our models. Furthermore, we sequentially log-transform and add additional control variables for robustness. In particular, the amount of money laundering in a country could, in and of itself, drive both STRs and convictions. Imagine, for example, that banks consistently file STRs on 10% of all money launderers and that 5% of all launderers are convicted regardless of the number of STRs filed. In this case, an increase in launderers “mechanically” leads to more STRs and convictions, creating a classic case of omitted variable bias (i.e., the illusion of a causal relationship). To address this, we would like to control for the amount of money laundering in a country; something we, for obvious reasons, do not have data on. Instead, we add $\log_{10}(SGDP_{(c,t)})$ as a control variable. This is motivated by the idea that $SGDP_{(c,t)}$, estimating the size of a country’s shadow economy, is a reasonable proxy for money laundering activities when we consider otherwise comparable countries. Indeed, we believe a larger shadow economy should increase demand for money laundering. The covert nature of money laundering means that we cannot test this idea (if we knew the amount of money laundering in a country, we would not need a proxy for it). Note, however, that we only use $\log_{10}(SGDP_{(c,t)})$ as a control variable. [^15] The reader may be concerned that $SGDP_{(c,t)}$ builds on an estimate from a model and is not directly observed. We agree with this concern but note that illicit financial flows (by their nature) cannot be observed. We also note that GDP figures, commonly used in econometric analysis, come from statistical models (though they might be very well-developed). We also add $\log_{10}(POL_{(c,t)})$ as a control variable. This is motivated by the idea that convictions for money laundering may depend on the number of police officers in a country. Indeed, in the classic Cobb-Douglas setup, one might think of $STR_{(c,t)}$ as capital and $POL_{(c,t)}$ as labor. Finally, we also try to control for time using, respectively, (i) $t$ (i.e., time) as a regressor (assuming a log-linear time trend) and (ii) time dummy variables. [^16] Our model in (i) may, assuming no other control variables, be specified as $\log_{10}(CON_{(c,t+j)})=\alpha+\beta\times\log_{10}(STR_{(c,t)})+\delta\times t$ where $\delta\in\mathbb{R}$ is a regression coefficient. Similarly, our model in (ii) may be specified as $\log_{10}(CON_{(c,t+j)})=\alpha+\beta\times\log_{10}(STR_{(c,t)})+\sum_{k=1}^{K}\delta_{(k)}\mathbf{1}_{\{t=k\}}$ where $K$ denotes the last year $t$ for which we have an observation of $STR_{(c,t)}$ to model and $\delta_{(1)},\dots,\delta_{(K)}$ are regression coefficients. For numerical stability, $t$ is measured relative to 2006. We dynamically drop observations if one of our control variables is missing. Standard errors are clustered at the unit (i.e., country) level as implemented by statsmodels [64].
@@ 5 Results
We first consider results from our pooled regression models, treating all observations as independent. Next, we consider results from our fixed effects regression models, allowing more robust inference by controlling for unobserved country-specific confounders. Subsection 5.1 presents our pooled regression results. Subsection 5.2 presents our fixed effects regression results. Throughout both subsections, we test the robustness of our results by varying dependent variables, sequentially excluding countries, and including control variables and time effects. We generally round results to two decimals and evaluate statistical significance at a $0.05$ $p$-level.
@@ 5.1 Pooled Regression Results
Table 3 presents our pooled regression results, considering multiple model specifications, numbered P1 to P4, and multiple dependent variables as explained in Section 4. In model P1, we only use $\log_{10}(STR_{(c,t)})$ as a regressor. In model P2, we include $\log_{10}(POP_{(c,t)})$ as a control variable. In model P3, we include $\log_{10}(SGDP_{(c,t)})$ as a control variable. Finally, in model P4, we include $\log_{10}(POL_{(c,t)})$ as a control variable. For illustration purposes, Figure 1 depicts model P1.
@@
Dependent variable: $\log_{10}(CON_{\left(c,t+1:3\right)})$
@@
Dependent variable: $\log_{10}(CON_{\left(c,t+1:5\right)})$
@@
Table 3: Pooled regression results. The table contains results for multiple model specifications, numbered P1 to P4, and multiple dependent variables as explained in Section 4. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level; and *** at the $0.001$ $p$-level.
@@
P1T 113 0.21 -0.87** (0.29) 0.36** (0.13) 0.04* (0.02)
P2T 113 0.33 -3.77** (1.41) 0.37** (0.12) 0.41 (0.21) 0.04* (0.02)
P3T 113 0.51 -8.62*** (2.06) 0.26 (0.13) 0.37 (0.20) 1.43** (0.49) 0.04* (0.02)
P4T 111 0.57 -12.04*** (1.98) 0.39** (0.13) 0.32* (0.16) 1.52*** (0.37) 1.28* (0.60) 0.04* (0.02)
Dependent variable: $\log_{10}(CON_{\left(c,t+1:3\right)})$
P1T 122 0.39 -1.48*** (0.24) 0.44*** (0.12) 0.09** (0.03)
P2T 122 0.52 -4.84*** (1.26) 0.45*** (0.12) 0.48* (0.19) 0.08** (0.03)
P3T 122 0.58 -8.33*** (2.00) 0.36** (0.13) 0.48** (0.17) 0.99** (0.38) 0.08** (0.03)
P4T 116 0.57 -9.53*** (2.68) 0.39* (0.16) 0.48** (0.17) 1.05** (0.40) 0.37 (0.68) 0.08** (0.03)
Dependent variable: $\log_{10}(CON_{\left(c,t+1:5\right)})$
P1T 80 0.47 -1.28*** (0.19) 0.43*** (0.08) 0.08* (0.03)
P2T 80 0.57 -3.84*** (1.10) 0.43*** (0.08) 0.37* (0.16) 0.08** (0.03)
P3T 80 0.68 -7.91*** (1.71) 0.34*** (0.10) 0.37** (0.14) 1.13*** (0.30) 0.08** (0.03)
P4T 80 0.70 -10.33*** (2.02) 0.42*** (0.12) 0.38** (0.13) 1.26*** (0.25) 0.74 (0.50) 0.06* (0.03)

### v2

@@ Abstract
Our results have important implications for international anti-money laundering efforts and policies. At best, the number of suspicious transaction reports and convictions for money laundering in a country appear to follow a sub-linear power law. At worst, there appears to be no or only a very weak (and insignificant) relationship between the number of suspicious transaction reports and convictions for money laundering in a country. Either way, our results suggest that financial regulators and supervisors should focus less on the number of suspicious transaction reports filed in a given country. Instead, they should focus more on how reports are utilized differently across countries and over time.
@@ 1 Introduction
A meta-analysis published by the United Nations Office on Drugs and Crime [61] suggests that money laundering amounts to 2.7% of the world’s combined gross domestic product (GDP). At the same time, scholars have long questioned the effectiveness of global anti-money laundering (AML) efforts [63, 58, 56, 62]. Such efforts are guided by the Financial Action Task Force (FATF), an intergovernmental organization founded in 1989. Among other things, FATF recommends that entities such as banks should be required to report suspicious transactions to national financial intelligence units [50]. The reports are known as suspicious transaction or activity reports (we use the former term) and are intended to help authorities detect and prosecute money laundering. In this paper, we investigate the relationship between suspicious transaction reports (STRs) and convictions for money laundering in the European Union (EU). Following the EU’s 4th AML directive, Member States must have “effective” AML systems and, among other things, require obligated entities to file STRs [45]. More or less implicitly, the EU directive and FATF’s AML recommendations thus assume that STRs are an effective tool to combat money laundering. This is also evident from the fact that STR counts are used as a measure of success in international AML evaluations [49, 48]. Our article explores this idea, asking if “more STRs lead to more convictions for money laundering?”
@@
Our results have important implications for international AML efforts and policies. At best, the number of STRs and convictions for money laundering in a country appears to follow a sub-linear power law. At worst, there appears to be no or only a very weak (and insignificant) relationship between the number of STRs and convictions for money laundering in a country. This would be a cause for serious concern about the effectiveness of current AML efforts and policies. Either way, our results suggest that financial regulators and supervisors should focus less on increasing the number of STRs filed in a given country. Instead, they should focus more on how STRs are utilized differently across countries and over time.
@@ 2 Related Literature
Scholars have long questioned the effectiveness of global AML efforts and policies [63, 58, 56, 62]. [62] calls AML “the world’s least effective policy experiment,” arguing that less than 0.1% of criminal proceeds are seized while AML costs vastly exceed assets recovered. The author partly ascribes the failure of the global AML system to an excessive focus on activity metrics (e.g., the number of STRs filed) over outcomes (e.g., the number of convictions achieved or reduced crime rates). Somewhat more diplomatic, [58] argue that evaluation of AML efforts is hindered by an absence of reliable data. Indeed, the authors note that AML evaluations and assessments (including those by FATF) mostly rely on rhetorical arguments and narrative judgments, providing little to no evidence on effectiveness. This fits with a broader pattern, noted by [57], whereby “governments and intergovernmental organizations have spent almost nothing on public research on these issues, while regular costly (…) international meetings service the global fight against this ill-understood phenomenon and to ‘evaluate’ these efforts.” In an extensive essay and review, [56] note that “the universal message from the modest research that has been conducted is that the proportion of SARs [a version of STRs] in high-reporting jurisdictions that are actually seriously followed up is low, though the extent to which this is inherent or merely resource-constrained remains unclear.” In a broader perspective, [56] note that “how well the [AML] system works in suppressing crimes and preventing terrorist acts is entirely a matter of speculation.”
@@ 3.3 Estimates of the Shadow Economy from the International Monetary Fund
We stress that “the shadow economy” and “money laundering” are distinct but related concepts. In particular, the shadow economy is defined as economic activity hidden from authorities [55, 54, 53, 69]. This will largely include activities that, in and of themselves, are legal. Furthermore, such activities do not necessarily involve money laundering; just as all criminal activities do not require money laundering. [^4] Imagine, for example, a day laborer who is paid in cash. The laborer’s work may, in and of itself, be legal. If the laborer fails to declare their income, it constitutes tax evasion. However, if the laborer simply uses their cash salary (including what should have been paid in taxes) to pay for daily expenses such as groceries, they may not be involved in any kind of money laundering.
@@
We use the European Sourcebook of Crime and Criminal Justice Statistics [37, 36, 38] as our source of police and conviction statistics. [^5] We extract and use raw statistics from the Sourcebook’s databases, available on the Sourcebook’s website [37]. These statistics may at times differ from those published in the print versions of the Sourcebook. The Sourcebook collects crime and criminal justice statistics from national correspondents. The 6th edition of the Sourcebook, for example, contains statistics from the Danish Ministry of Justice, the Latvian Statistical Bureau, and numerous academics from across Europe. Notably, the Sourcebook aims to collect standardized statistics; a prerequisite for comparing statistics across countries. [^6] As an alternative data source, we considered using mutual evaluation reports from FATF. However, statistics in these are not standardized. For example, [47] combines statistics on money laundering and stolen goods. While the Sourcebook aims to collect standardized statistics, these may deviate. In this work, we simply use the statistics as given in the Sourcebook. The Sourcebook explicitly defines money laundering as “specific financial transactions to conceal the identity, source, and/or destination of money or non-monetary property deriving from criminal activities.” We collect the following variables from the Sourcebook:
@@
Figures 2 and 3 illustrate the number of observations of $CON_{(c,t)}$ per country and year in our dataset. As can be seen, there are some countries for which we have no observations. Still, our dataset covers Northern, Southern, Eastern, and Western Europe. As can also be seen, the number of observations increases towards the end of our data collection period. We believe this is due to improvements in the way data was collected in the Sourcebook; not interaction effects between STRs and convictions. Interestingly, the Sourcebook also contains statistics on convictions for some predicate crimes to money laundering; we consider these in Appendix C.
@@ 4 Methodology
denote the average number of convictions in country $c$ per year up to and including $H$ years after $t$. We measure $CON_{(c,t+j)}$ and $CON_{(c,t+1:H)}$ per 100,000 capita using $POP_{(c,t)}$. [^11] This is based on the idea that $CON_{(c,t+j)}$ and $CON_{(c,t+1:H)}$ are determined by $STR_{(c,t)}$, i.e., the number of STRs in year $t$, not years t+1,,t+Ht+1,...,t+H. For robustness, we also tried measuring $CON_{(c,t+j)}$ and $CON_{(c,t+1:H)}$ using population counts in year $t+j$ and the average population count over years t+1,,t+Ht+1,...,t+H. This does not materially change our conclusions. We now consider and model:
@@ 4.1 Pooled Regression
where $A\in\mathbb{R}_{>0}$ measures a “scale of production” and $\beta\in\mathbb{R}$ measures how convictions respond to a change in STRs. [^13] As an alternative to the power-law formulation, we also considered a simple linear formulation; see Appendix B. To fit the model, we use a log-transformation, yielding an ordinary least squares regression
@@
Motivated by the observation that we have per capita ratios on both sides of equation (3) (note that both $CON_{(c,t+j)}$ and $STR_{(c,t)}$ are measured per 100,000 capita) and that a country’s population size might have a stand-alone effect on conviction rates, we include $log_{10}(POP_{(c,t)})$ as a control variable in the model in (3). [^15] Algebraically, $\log_{10}(y/n)=\alpha+\beta\,\log_{10}(x/n)+u\iff\log_{10}y=\alpha+\beta\,\log_{10}(x)+(1-\beta)\,\log_{10}(n)+u$. Thus, omitting $\log_{10}(n)$ (or, in our case, $\log_{10}(POP_{(c,t)})$) as a control variable imposes a (hidden) restriction that its coefficient must equal $1-\beta$. Note, also, that measuring variables per (100,000) capita shifts the intercept of our models. Furthermore, we sequentially log-transform and add additional control variables for robustness. In particular, the amount of money laundering in a country could, in and of itself, drive both STRs and convictions. Imagine, for example, that banks consistently file STRs on 10% of all money launderers and that 5% of all launderers are convicted regardless of the number of STRs filed. In this case, an increase in launderers “mechanically” leads to more STRs and convictions, creating a classic case of omitted variable bias (i.e., the illusion of a causal relationship). To address this, we would like to control for the amount of money laundering in a country; something we, for obvious reasons, do not have data on. Instead, we add $\log_{10}(SGDP_{(c,t)})$ as a control variable. This is motivated by the idea that $SGDP_{(c,t)}$, estimating the size of a country’s shadow economy, is a reasonable proxy for money laundering activities when we consider comparable countries. Indeed, we find it is reasonable to suspect that a larger shadow economy, all else being equal, might increase demand for money laundering. Furthermore, a larger shadow economy might also, in itself, influence the number of STRs being filed. The covert nature of money laundering means that we cannot directly investigate the relationship between a country’s shadow economy and the amount of money laundering in said country (if we knew the amount of money laundering in a country, we would not need a proxy for it). Note, however, that we only use $\log_{10}(SGDP_{(c,t)})$ as a control variable. [^16] The reader may be concerned that $SGDP_{(c,t)}$ builds on an estimate from a model and is not directly observed. We agree with this concern but note that illicit financial flows (by their nature) cannot be observed. We also note that GDP figures, commonly used in econometric analysis, come from statistical models (though they might be very well-developed). We also add $\log_{10}(POL_{(c,t)})$ as a control variable. This is motivated by the idea that convictions for money laundering may depend on the number of police officers in a country. Indeed, in the classic Cobb-Douglas setup, one might think of $STR_{(c,t)}$ as capital and $POL_{(c,t)}$ as labor. Finally, we also try to control for time using, respectively, (i) $t$ (i.e., time) as a regressor (assuming a log-linear time trend) and (ii) time dummy variables. [^17] Our model in (i) may, assuming no other control variables, be specified as $\log_{10}(CON_{(c,t+j)})=\alpha+\beta\times\log_{10}(STR_{(c,t)})+\delta\times t$ where $\delta\in\mathbb{R}$ is a regression coefficient. Similarly, our model in (ii) may be specified as $\log_{10}(CON_{(c,t+j)})=\alpha+\beta\times\log_{10}(STR_{(c,t)})+\sum_{k=1}^{K}\delta_{(k)}\mathbf{1}_{\{t=k\}}$ where $K$ denotes the last year $t$ for which we have an observation of $STR_{(c,t)}$ to model and $\delta_{(1)},\dots,\delta_{(K)}$ are regression coefficients. For numerical stability, $t$ is measured relative to 2006. We dynamically drop observations if one of our control variables is missing. Standard errors are clustered at the unit (i.e., country) level as implemented by statsmodels [64].
@@ 5 Results
We first consider results from our pooled regression models, treating all observations as independent. Next, we consider results from our fixed effects regression models, allowing more robust inference by controlling for unobserved country-specific confounders. Subsection 5.1 presents our pooled regression results. Subsection 5.2 presents our fixed effects regression results. Throughout both subsections, we test the robustness of our results by varying dependent variables, sequentially excluding countries, and including control variables and time effects. We generally round results to two decimals and evaluate statistical significance at a $0.05$ $p$-level. The European Sourcebook of Crime and Criminal Justice Statistics also contains statistics on convictions for some predicate crimes to money laundering: we present results using these in Appendix C.
@@ 5.1 Pooled Regression Results
Table 3 presents our pooled regression results, considering multiple model specifications, named P1 to P4, and multiple dependent variables as explained in Section 4. In model P1, we only use $\log_{10}(STR_{(c,t)})$ as a regressor. In model P2, we include $\log_{10}(POP_{(c,t)})$ as a control variable. In model P3, we include $\log_{10}(SGDP_{(c,t)})$ as a control variable. Finally, in model P4, we include $\log_{10}(POL_{(c,t)})$ as a control variable. For illustration purposes, Figure 1 depicts model P1. A Q-Q plot of residuals is displayed in Appendix A.
@@
Dependent variable: $\log_{10}\left(CON_{(c,t+1:3)}\right)$
@@
Dependent variable: $\log_{10}\left(CON_{(c,t+1:5)}\right)$
@@
Table 3: Pooled regression results. The table contains results for multiple model specifications, named P1 to P4, and multiple dependent variables as explained in Section 4. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level; and *** at the $0.001$ $p$-level.
@@
PT1 113 0.21 -0.87** (0.29) 0.36** (0.13) 0.04* (0.02)
PT2 113 0.33 -3.77** (1.41) 0.37** (0.12) 0.41 (0.21) 0.04* (0.02)
PT3 113 0.51 -8.62*** (2.06) 0.26 (0.13) 0.37 (0.20) 1.43** (0.49) 0.04* (0.02)
PT4 111 0.57 -12.04*** (1.98) 0.39** (0.13) 0.32* (0.16) 1.52*** (0.37) 1.28* (0.60) 0.04* (0.02)
Dependent variable: $\log_{10}\left(CON_{(c,t+1:3)}\right)$
PT1 122 0.39 -1.48*** (0.24) 0.44*** (0.12) 0.09** (0.03)
PT2 122 0.52 -4.84*** (1.26) 0.45*** (0.12) 0.48* (0.19) 0.08** (0.03)
PT3 122 0.58 -8.33*** (2.00) 0.36** (0.13) 0.48** (0.17) 0.99** (0.38) 0.08** (0.03)
PT4 116 0.57 -9.53*** (2.68) 0.39* (0.16) 0.48** (0.17) 1.05** (0.40) 0.37 (0.68) 0.08** (0.03)
Dependent variable: $\log_{10}\left(CON_{(c,t+1:5)}\right)$
PT1 80 0.47 -1.28*** (0.19) 0.43*** (0.08) 0.08* (0.03)
PT2 80 0.57 -3.84*** (1.10) 0.43*** (0.08) 0.37* (0.16) 0.08** (0.03)
PT3 80 0.68 -7.91*** (1.71) 0.34*** (0.10) 0.37** (0.14) 1.13*** (0.30) 0.08** (0.03)
PT4 80 0.70 -10.33*** (2.02) 0.42*** (0.12) 0.38** (0.13) 1.26*** (0.25) 0.74 (0.50) 0.06* (0.03)